# Contemporary Recommendation Systems on Big Data and Their Applications: A Survey

### 08 Mar 2024 v3

@@ Abstract
This survey paper conducts a comprehensive analysis of the evolution and contemporary landscape of recommendation systems, which have been extensively incorporated across a myriad of web applications. It delves into the progression of personalized recommendation methodologies tailored for online products or services, organizing the array of recommendation techniques into four main categories: content-based, collaborative filtering, knowledge-based, and hybrid approaches, each designed to cater to specific contexts. The document provides an in-depth review of both the historical underpinnings and the cutting-edge innovations in the domain of recommendation systems, with a special focus on implementations leveraging big data analytics. It further outlines and explores the predominant challenges encountered in the current generation of recommendation systems, including issues related to data sparsity, scalability, and the imperative for diversified recommendation outputs. The survey underscores these challenges as promising directions for subsequent research endeavors within the discipline. Additionally, the paper examines various real-life applications driven by recommendation systems, addressing the hurdles involved in seamlessly integrating these systems into everyday life. Ultimately, the survey underscores how the advancements in recommendation systems, propelled by big data technologies, have the potential to significantly enhance real-world experiences.
@@
Comprehensive Classification and Analysis of Recommendation Techniques: The paper meticulously categorizes the myriad of recommendation system methodologies into four principal types: content-based, collaborative filtering, knowledge-based, and hybrid approaches. It offers a detailed exploration of how these methodologies have evolved over time and their application in providing personalized recommendations across various online platforms. This classification serves as a foundational framework for understanding the diverse strategies employed in tailoring user experiences and addresses the unique challenges and requirements specific to each approach.
Examination of Challenges and Future Research Directions: Highlighting the current challenges faced by recommendation systems, such as data sparsity, scalability, and the need for diversifying recommendation outputs, the paper not only provides a critical analysis of these issues but also positions them as promising avenues for future research. This focus underscores the dynamic nature of the field and encourages the exploration of innovative solutions to enhance the effectiveness and efficiency of recommendation systems in handling large-scale data and meeting the evolving preferences of users.
Real-World Applications and Impact on Society: The survey extends its analysis beyond theoretical models to examine the practical implementation and impact of recommendation systems in everyday life. It explores their application across various domains including marketing, healthcare, and governance, and discusses how they influence consumer behavior and societal trends. By highlighting the potential of recommendation systems to promote sustainable lifestyles and improve decision-making processes, the paper sheds light on the significant role these technologies play in shaping our digital and real-world experiences.
@@ 3.2 Modern Recommendation Systems based on the Big Data
4 Recommendation System Application
@@ 4.4 Recommendation System in Healthcare
Clinical Decision Support. One of the primary applications of recommendation systems in healthcare is in clinical decision support. These systems analyze electronic health records (EHRs), medical literature, and patient data to assist healthcare professionals in making informed decisions about diagnostics, treatment plans, and interventions. By providing relevant and evidence-based information, recommendation systems contribute to more accurate and personalized patient care, potentially reducing diagnostic errors and improving overall healthcare outcomes bates2003impact ; shojania2009effects .
Patient-Centered Care. In the era of patient-centered care, recommendation systems play a crucial role in tailoring healthcare services to individual patient needs. These systems analyze patient preferences, demographics, and health histories to generate personalized recommendations for treatment options, preventive measures, and lifestyle modifications. By fostering patient engagement and empowerment, recommendation systems contribute to a more collaborative and effective healthcare relationship between providers and patients coulter2007effectiveness ; street2009communication .
Resource Optimization. Recommendation systems help optimize healthcare resources by streamlining processes and improving operational efficiency. For instance, in hospital management, these systems can suggest optimal bed allocation, appointment scheduling, and resource utilization based on historical data and real-time information. By minimizing bottlenecks and enhancing resource allocation, recommendation systems contribute to cost-effectiveness and improved service delivery li2017survey ; brailsford2009comprehensive .
Telehealth and Remote Monitoring. With the rise of telehealth and remote monitoring, recommendation systems support healthcare providers in delivering virtual care. These systems analyze patient-generated health data from wearable devices, monitoring tools, and telehealth platforms to provide timely recommendations for interventions, medication adjustments, or lifestyle modifications. This real-time support contributes to proactive healthcare management, particularly for patients with chronic conditions fatehi2018clinician ; dorsey2016state .
Collaborative Healthcare Networks. Recommendation systems facilitate collaboration and knowledge sharing among healthcare professionals through the creation of collaborative networks. By analyzing expertise, research interests, and clinical experiences, these systems connect healthcare professionals for consultations, research collaborations, and second opinions. This fosters a culture of continuous learning and knowledge dissemination within the healthcare community alshammary2019collaborative ; choi2018systematic .
@@ 5 Datasets for Recommendation Systems
Datasets play a pivotal role in the development, evaluation, and benchmarking of recommendation systems. The choice of dataset depends on factors such as the application domain, the type of recommendation task, and the specific research objectives. In this section, we provide a summary of some commonly used datasets in recommendation systems research:
@@ 5.1 MovieLens
The MovieLens dataset is one of the most widely used datasets for collaborative filtering-based recommendation systems. It contains movie ratings provided by users, along with movie metadata such as titles, genres, and release years. MovieLens datasets are available in various sizes, ranging from small-scale datasets suitable for initial experimentation to large-scale datasets for comprehensive evaluations.
@@ 5.2 Amazon Reviews
Amazon provides extensive datasets comprising user reviews and ratings for a wide range of products available on its platform, including books, electronics, clothing, and more. These datasets offer rich information about user preferences, product attributes, and user-item interactions, making them valuable resources for research in recommendation systems, sentiment analysis, and e-commerce analytics.
@@ 5.3 Netflix Prize Dataset
The Netflix Prize dataset is a large-scale dataset of movie ratings collected from Netflix users. It was released as part of the Netflix Prize competition, a renowned competition aimed at improving the performance of recommendation algorithms. The dataset contains millions of ratings provided by users over several years, along with additional contextual information such as user demographics and temporal dynamics.
@@
The Last.fm dataset comprises music listening histories of users on the Last.fm platform, including details about artists, albums, and user preferences. It offers insights into user behavior and preferences in the context of music consumption, making it valuable for research in music recommendation systems, personalized playlists, and music discovery applications.
@@ 5.5 Yelp Dataset
The Yelp dataset consists of user reviews and ratings for businesses across various categories, including restaurants, hotels, and local services. It provides rich information about user opinions, business attributes, and geographic locations, enabling research in recommendation systems, sentiment analysis, and location-based services.
@@ 6.1 Interpretable Recommender Systems
Research could focus on developing recommendation models that not only provide accurate predictions but also offer transparent explanations for their recommendations. Techniques such as model-agnostic explanation methods, rule-based systems, or attention mechanisms can be explored to enhance interpretability.
@@ 6.2 Fairness and Bias
Future research might concentrate on devising fairness-aware recommendation algorithms that mitigate bias and ensure equitable treatment across diverse user groups. This could involve developing fairness metrics, fairness-aware loss functions, or debiasing techniques tailored for recommendation systems.
@@ 6.3 Context-Aware Recommendations
Advanced research may explore leveraging deep learning architectures, such as recurrent neural networks (RNNs) or transformers, to effectively capture and utilize contextual information in recommendation models. Additionally, techniques like multi-modal learning could be investigated to incorporate various types of context, such as textual, visual, or temporal cues.
@@ 6.4 Cold-Start Problem
Novel research directions could involve exploring transfer learning approaches, where knowledge learned from related domains or auxiliary data sources is transferred to alleviate the cold-start problem. Hybrid recommendation techniques combining collaborative filtering, content-based methods, and knowledge graphs could also be investigated to handle cold-start scenarios more effectively.
@@ 6.5 Long-Tail Recommendations
Research efforts could focus on developing specialized algorithms for long-tail recommendations, such as mixture models, probabilistic graphical models, or ensemble methods tailored to capture rare item preferences. Techniques like active learning or diversity-promoting algorithms could also be explored to enhance the coverage of long-tail items in recommendations.
@@ 6.6 Dynamic and Adaptive Recommendations
Future research might investigate reinforcement learning frameworks for recommendation systems, where agents learn optimal recommendation policies through interaction with users and feedback. Additionally, online learning algorithms capable of continuously updating recommendation models in response to evolving user preferences and feedback could be explored.
@@ 6.7 Privacy-Preserving Recommendations
Research could explore techniques such as federated learning, where recommendation models are trained across decentralized data sources without sharing raw user data. Differential privacy mechanisms could also be integrated into recommendation algorithms to ensure individual user privacy while maintaining recommendation utility.
@@ 6.8 Multi-Stakeholder Recommendations
Advanced research may focus on developing multi-objective optimization frameworks for recommendation systems, considering the diverse interests of users, providers, and advertisers simultaneously. Game-theoretic approaches could be employed to model interactions among stakeholders and design recommendation strategies that optimize collective utility while addressing conflicting objectives.
@@ 7 Summary
The manuscript delves into a comprehensive historical review and a critical examination of the state-of-the-art methodologies in recommendation systems, with a special emphasis on the pioneering developments brought about by the advent of big data analytics. Additionally, this paper scrutinizes the prevalent challenges encountered in contemporary recommendation systems, such as data sparsity, scalability issues, and the need for diversity in recommendations, proposing these hurdles as fertile grounds for future research endeavors.

### 13 Mar 2024 v4

@@ Abstract
This survey paper conducts a comprehensive analysis of the evolution and contemporary landscape of recommendation systems, which have been extensively incorporated across a myriad of web applications. It delves into the progression of personalized recommendation methodologies tailored for online products or services, organizing the array of recommendation techniques into four main categories: content-based, collaborative filtering, knowledge-based, and hybrid approaches, each designed to cater to specific contexts. The document provides an in-depth review of both the historical underpinnings and the cutting-edge innovations in the domain of recommendation systems, with a special focus on implementations leveraging big data analytics. The paper also highlights the utilization of prominent datasets such as MovieLens, Amazon Reviews, Netflix Prize, Last.fm, and Yelp in evaluating recommendation algorithms. It further outlines and explores the predominant challenges encountered in the current generation of recommendation systems, including issues related to data sparsity, scalability, and the imperative for diversified recommendation outputs. The survey underscores these challenges as promising directions for subsequent research endeavors within the discipline. Additionally, the paper examines various real-life applications driven by recommendation systems, addressing the hurdles involved in seamlessly integrating these systems into everyday life. Ultimately, the survey underscores how the advancements in recommendation systems, propelled by big data technologies, have the potential to significantly enhance real-world experiences.
@@
Comprehensive Analysis and Methodological Classification: It provides a thorough exploration of the evolution and current landscape of recommendation systems, categorizing the techniques into content-based, collaborative filtering, knowledge-based, and hybrid approaches. This classification is designed to address the unique requirements of various online products and services, emphasizing the role of big data analytics in driving innovations within this domain.
Evaluation Using Prominent Datasets and Challenges: The paper emphasizes the use of key datasets, such as MovieLens, Amazon Reviews, Netflix Prize, Last.fm, and Yelp, for evaluating the performance of recommendation algorithms. It also outlines significant challenges faced by modern recommendation systems, including data sparsity, scalability issues, and the need for diversified recommendations, positioning these challenges as avenues for future research.
Real-life Applications and Impact on Society: By examining the integration of recommendation systems in various domains such as marketing, governance, and healthcare, the paper underscores their significance in enhancing real-world experiences. It highlights how advancements in recommendation technologies, propelled by big data, are shaping user experiences and influencing societal trends, offering insights into their potential for future societal impact.
@@ 3.2 Modern Recommendation Systems based on the Big Data
4 Applications of Modern Recommendation System
@@ 4.4 Recommendation System in Healthcare
4.4.1 Clinical Decision Support
. One of the primary applications of recommendation systems in healthcare is in clinical decision support. These systems analyze electronic health records (EHRs), medical literature, and patient data to assist healthcare professionals in making informed decisions about diagnostics, treatment plans, and interventions. By providing relevant and evidence-based information, recommendation systems contribute to more accurate and personalized patient care, potentially reducing diagnostic errors and improving overall healthcare outcomes bates2003impact ; shojania2009effects .
4.4.2 Patient-Centered Care
. In the era of patient-centered care, recommendation systems play a crucial role in tailoring healthcare services to individual patient needs. These systems analyze patient preferences, demographics, and health histories to generate personalized recommendations for treatment options, preventive measures, and lifestyle modifications. By fostering patient engagement and empowerment, recommendation systems contribute to a more collaborative and effective healthcare relationship between providers and patients coulter2007effectiveness ; street2009communication .
4.4.3 Resource Optimization
. Recommendation systems help optimize healthcare resources by streamlining processes and improving operational efficiency. For instance, in hospital management, these systems can suggest optimal bed allocation, appointment scheduling, and resource utilization based on historical data and real-time information. By minimizing bottlenecks and enhancing resource allocation, recommendation systems contribute to cost-effectiveness and improved service delivery li2017survey ; brailsford2009comprehensive .
4.4.4 Telehealth and Remote Monitoring
. With the rise of telehealth and remote monitoring, recommendation systems support healthcare providers in delivering virtual care. These systems analyze patient-generated health data from wearable devices, monitoring tools, and telehealth platforms to provide timely recommendations for interventions, medication adjustments, or lifestyle modifications. This real-time support contributes to proactive healthcare management, particularly for patients with chronic conditions fatehi2018clinician ; dorsey2016state .
4.4.5 Collaborative Healthcare Networks
. Recommendation systems facilitate collaboration and knowledge sharing among healthcare professionals through the creation of collaborative networks. By analyzing expertise, research interests, and clinical experiences, these systems connect healthcare professionals for consultations, research collaborations, and second opinions. This fosters a culture of continuous learning and knowledge dissemination within the healthcare community alshammary2019collaborative ; choi2018systematic .
@@ 5 Datasets for Recommendation Systems
Datasets play a pivotal role in the development, evaluation, and benchmarking of recommendation systems. The choice of dataset depends on factors such as the application domain, the type of recommendation task, and the specific research objectives. In this section, we provide a summary of some commonly used datasets, as shown in Table. 2, in recommendation systems research:
Dataset Top Models Metrics
Extra
Training
Data
Year
MovieLens 100K GHRS GHRS 0.887(RSME) N 2021
GLocal-K GlocalK 0.889(RSME) N 2021
MG-GAT MGGAT 0.890(RSME) Y 2020
MovieLens 1M GLocal-K GlocalK 0.823(RSME) N 2021
Sparse FC SparseFC 0.824(RSME) N 2018
CF-NADE CFNADE 0.829(RSME) N 2016
MovieLens 10M
Bayesian
timeSVD++
flipped
 BayesiantimeSVD++
0.749(RSME) N 2019
Bayesian
timeSVD++
 BayesiantimeSVD++
0.752(RSME) N 2019
Bayesian
SVD++
 BayesiantimeSVD++
0.756(RSME) N 2019
MovieLens 20M VASP VASP 0.448(nDCG@100) Y 2021
H+Vamp
Gated
 H+VAMP
0.445(nDCG@100) Y 2019
RecVAE Recvae 0.442(nDCG@100) N 2019
Amazon-Book SSCF SSCF 0.065(nDCG@20) N 2022
SANSA SANSA 0.064(nDCG@20) N 2023
BSPM-LM BSPMLM 0.061(nDCG@20) N 2022
Netflix
H+Vamp
Gated
 H+VAMP
0.409(nDCG@100) N 2019
RecVAE Recvae 0.394(nDCG@100) N 2019
EASE EASE 0.393(nDCG@100) N 2019
Last.FM Ekar EKAR 0.248(HR@10) N 2019
HAKG HAKG 0.093(nDCG@20) N 2022
KGNN-LS KGNNLS 0.370(Recall@100) N 2019
Yelp DGRec DGRec 0.1427(nDCG) N 2019
Gowalla BSPM-EM BSPM-EM 0.160(nDCG@20) N 2022
MGDCF Mgdcf 0.159(nDCG@20) N 2024
UltraGCN UltraGCN 0.158(nDCG@20) N 2021
Table 2: Datasets and compared methods for recommendation systems
@@ 5.1 MovieLens
The MovieLens dataset is one of the most widely used datasets for collaborative filtering-based recommendation systems. It contains movie ratings provided by users, along with movie metadata such as titles, genres, and release years movielens . MovieLens datasets are available in various sizes (e.g. 100K, 1M, 10M, 20M) ranging from small-scale datasets suitable for initial experimentation to large-scale datasets for comprehensive evaluations.
@@ 5.2 Amazon Reviews
Amazon provides extensive datasets comprising user reviews and ratings for a wide range of products available on its platform, including books, electronics, clothing, and more amazon_reviews . These datasets offer rich information about user preferences, product attributes, and user-item interactions, making them valuable resources for research in recommendation systems, sentiment analysis, and e-commerce analytics.
@@ 5.3 Netflix Prize Dataset
The Netflix Prize dataset is a large-scale dataset of movie ratings collected from Netflix users. It was released as part of the Netflix Prize competition, a renowned competition aimed at improving the performance of recommendation algorithms netflix_prize . The dataset contains millions of ratings provided by users over several years, along with additional contextual information such as user demographics and temporal dynamics.
@@
The Last.fm dataset comprises music listening histories of users on the Last.fm platform, including details about artists, albums, and user preferences lastfm . It offers insights into user behavior and preferences in the context of music consumption, making it valuable for research in music recommendation systems, personalized playlists, and music discovery applications.
@@ 5.5 Yelp Dataset
The Yelp dataset consists of user reviews and ratings for businesses across various categories, including restaurants, hotels, and local services yelp . It provides rich information about user opinions, business attributes, and geographic locations, enabling research in recommendation systems, sentiment analysis, and location-based services.
@@ 6.1 Interpretable Recommender Systems
Research could focus on developing recommendation models that not only provide accurate predictions but also offer transparent explanations for their recommendations explanation_methods . Techniques such as model-agnostic explanation methods, rule-based systems, or attention mechanisms can be explored to enhance interpretability.
@@ 6.2 Fairness and Bias
Future research might concentrate on devising fairness-aware recommendation algorithms that mitigate bias and ensure equitable treatment across diverse user groups fairness_recsys . This could involve developing fairness metrics, fairness-aware loss functions, or debiasing techniques tailored for recommendation systems.
@@ 6.3 Context-Aware Recommendations
Advanced research may explore leveraging deep learning architectures, such as recurrent neural networks (RNNs) or transformers, to effectively capture and utilize contextual information in recommendation models context_aware_recsys . Additionally, techniques like multi-modal learning could be investigated to incorporate various types of context, such as textual, visual, or temporal cues.
@@ 6.4 Cold-Start Problem
Novel research directions could involve exploring transfer learning approaches, where knowledge learned from related domains or auxiliary data sources is transferred to alleviate the cold-start problem transfer_learning_recsys . Hybrid recommendation techniques combining collaborative filtering, content-based methods, and knowledge graphs could also be investigated to handle cold-start scenarios more effectively.
@@ 6.5 Long-Tail Recommendations
Research efforts could focus on developing specialized algorithms for long-tail recommendations, such as mixture models, probabilistic graphical models, or ensemble methods tailored to capture rare item preferences long_tail_recsys . Techniques like active learning or diversity-promoting algorithms could also be explored to enhance the coverage of long-tail items in recommendations.
@@ 6.6 Dynamic and Adaptive Recommendations
Future research might investigate reinforcement learning frameworks for recommendation systems, where agents learn optimal recommendation policies through interaction with users and feedback reinforcement_learning_recsys . Additionally, online learning algorithms capable of continuously updating recommendation models in response to evolving user preferences and feedback could be explored.
@@ 6.7 Privacy-Preserving Recommendations
Research could explore techniques such as federated learning, where recommendation models are trained across decentralized data sources without sharing raw user data privacy_preserving_recsys . Differential privacy mechanisms could also be integrated into recommendation algorithms to ensure individual user privacy while maintaining recommendation utility.
@@ 6.8 Multi-Stakeholder Recommendations
Advanced research may focus on developing multi-objective optimization frameworks for recommendation systems, considering the diverse interests of users, providers, and advertisers simultaneously multi_stakeholder_recsys . Game-theoretic approaches could be employed to model interactions among stakeholders and design recommendation strategies that optimize collective utility while addressing conflicting objectives.
@@ 7 Summary
The manuscript delves into a comprehensive historical review and a critical examination of the state-of-the-art methodologies in recommendation systems, with a special emphasis on the pioneering developments brought about by the advent of big data analytics. Notably, this paper highlights the utilization of prominent datasets such as MovieLens, Amazon Reviews, Netflix Prize, Last.fm, and Yelp in evaluating recommendation algorithms. Additionally, this paper scrutinizes the prevalent challenges encountered in contemporary recommendation systems, such as data sparsity, scalability issues, and the need for diversity in recommendations, proposing these hurdles as fertile grounds for future research endeavors.